Two electrons are passing 27 mm apart. What is the electric repulsive force that they exert on each other? (The value of k is 9.0
Answer #1
1.2 × 1010 N
Answer #2
Repulsive Force = kq^2/r^2

= (9*10^9)*(1.6*10^-19)^2/(27*10^-3)^2

= 3.2 × 10-25 N
Answer #3

3.2 × 10-25 N